OpenStudy (anonymous):

-(4)^-2 A.-1/-2^4 B.4^2 C.1/4^2 D.-1/4^2

OpenStudy (hoblos):

in general.. a^-n = 1/a^n so -(4)^-2 = -1/4^2

Parth (parthkohli):

Use the property. \(\Large \color{MidnightBlue}{\Rightarrow x^{-n} = {1 \over x^n} }\)